# Demonstrator System Testing and Performance for the ATLAS ITk Pixel Detector for HL-LHC

### v1

@@ 1 Introduction
The integrated luminosity of the LHC will be increased tenfold over its original design threshold, for the transition into the High Luminosity-LHC (HL-LHC) [1]. This will significantly increase the LHC collision reach, leading to higher detector occupancy and accelerated radiation damage. To address these challenges, the current detectors must be replaced and upgraded to cope with the new requirements and demands of the HL-LHC, with faster readout detectors, and with higher granularity. Specifically, the ATLAS experiment will upgrade its Inner Detector (ID) to a new all-silicon Inner Tracker (ITk) system [2].
The ITk comprises of a four-layer strip detector with planar sensors and a five-layer hybrid pixel detector which is the focus of our paper, covering approximately 13m² and featuring around 10,000 hybrid modules. The inner layer are equipped with 3D sensors, while planar sensors are deployed on the outer layers. This new design promises enhanced performance but also presents complex operational demands, requiring extensive system testing and integration studies to mitigate the system complexities.
@@ 2 Demonstrator Setup and Associated Sub-systems
The demonstrator setup is built around two interconnected systems as figure 1 illustrates: on-detector and off-detector services. The on-detector services include module power supplies, a $CO_{2}$ evaporative cooling plant for efficient cooling and a dedicated Monitoring of Pixel System (MOPS) [3] for monitoring operational status, and for tracking individual power channels and module temperatures. Serial powering groups modules into chains powered by a constant current, with SLDO regulators for shunting excess current.
@@
The demonstrator’s testing program includes thermal and electrical tests, with system checks for MOPS, serial powering, and cooling efficiency. These tests identify issues in integration, module performance, and NTC consistency, ensuring the demonstrator’s readiness for large-scale production.
@@ 3 Readout Chain, DCS, and Data Transmission
The demonstrator system minimizes cabling needs through serial powering and uses electrical services to route data and command signals to Opto-converters. Signal aggregation and electrical-to-optical conversion occur in the Opto-box outside the test environment for radiation shielding and component accessibility. Each quad module uplink is supported by 1.28 Gb/s fibers to manage high data volumes, and cable lengths reflect the real setup for bandwidth and transmission reliability testing.
@@
Nevertheless, the DCS controls component power and integrates it into a finite state machine (FSM) interface, simplifying setup operation and safety by the interlock. The FSM allows for quick navigation between component states via a graphical user interface (GUI), enabling efficient operation for the ITk pixel detector’s future deployment.
@@ 4 Local Loaded Supports and $CO_{2}$ Cooling System
Precision in assembly is critical, especially regarding the layer of glue used to attach modules to the stave [9]. Excessive glue can shift module positions and reduce the thermal interface, negatively impacting heat dissipation from modules to the cooling system. Each stave integrates four key elements for optimized thermal management: thin-walled titanium tubes that circulate $CO_{2}$ coolant, base cooling blocks strategically placed to create thermally conductive interfaces, and carbon-fiber composites that form the primary support structure as figure 3 explicitly shows, ensuring stability while managing temperature fluctuations.
@@
Figure 3: Figure (left) shows $CO_{2}$ flow from inlet to exhaust with the module added to the cell-loading structure on the stave. Figure (right) shows thermal variations in NTC measurements were observed between single-module cooling jig tests and via the $CO_{2}$ cooling pipe.
@@ 5 RD53A Module Reliability in Full-System Tests
Figure 4: IHR stave installed inside the test box and mounted with 10 cell-loaded RD53A quad modules attached with pig-tail cables via the patch panel 0 (PP0).
A key objective of the OB demonstrator project is to validate the loading concepts crucial to ITk detector construction with the overall sub-systems described in section 3. Using detector prototypes equipped with RD53A electronics [10] as in Figure 4, the electrical scans taken at each stage of the production flow ensure the reliability of the module’s performance. A large set of electrical scans are needed to characterize the functionality of individual module components using standardized quality control (QC) measures across all production stages. To ensure direct inspections of any potential degradation by comparing the output to the previous and following stages. A final round of electrical scans evaluates the fully loaded modules in the ITk OB demonstrator setup, which is illustrated for 10 modules loaded on the IHR in Figure 5 of the OB demonstrator at SR1 in CERN. Which features the full-system test of DAQ, $CO_{2}$ cooling, SLDO serial powering chains, and collective module testing. The performance across all six testing stages demonstrates consistency, with minimal variation across stages despite environmental differences and fundamentally different setups.
Figure 5: The figure illustrates the multiple testing stages that RD53a modules undergo throughout various production steps. Upon arrival at CERN, an initial test is conducted to evaluate baseline functionality. Then, wire bonding is secured by coating for protection. The modules then proceed to cell loading on the back of the module as shown in Figure 3 and pig-tail assembly before entering the final OB demonstrator system test phase. Notably, noise scan performance remains consistent across all stages, indicating stable behavior for the IHR modules in the OB demonstrator..

### v2

@@ 1 Introduction
The integrated luminosity of the LHC will be increased tenfold over its original design threshold, for the transition into the High Luminosity-LHC (HL-LHC) [1]. This will significantly increase the LHC collision reach, leading to higher detector occupancy and accelerated radiation damage. To address these challenges, the current detectors must be replaced and upgraded to cope with the new requirements and demands of the HL-LHC, with faster readout detectors, and with higher granularity. Specifically, the ATLAS experiment [2] will upgrade its Inner Detector (ID) to a new all-silicon Inner Tracker (ITk) system [3].
The ITk comprises of a four-layer strip detector with planar sensors and a five-layer hybrid pixel detector which is the focus of our paper, covering approximately 13 m² and featuring around 10,000 hybrid modules. The inner layer is equipped with 3D sensors, while planar sensors are deployed on the outer layers. This new design promises enhanced performance but also presents complex operational demands, requiring extensive system testing and integration studies to mitigate the system complexities.
@@ 2 Demonstrator Setup and Associated Sub-systems
The demonstrator setup is built around two interconnected systems as Figure 1 illustrates: on-detector and off-detector services. The on-detector services include module power supplies, a $CO_{2}$ evaporative cooling plant for efficient cooling and a dedicated Monitoring of Pixel System (MOPS) [4] for monitoring operational status, and for tracking individual power channels and module temperatures. Serial powering groups modules into chains powered by a constant current, with Shunt Low Dropout (SLDO) regulators for shunting excess current.
@@
The demonstrator’s testing program includes thermal and electrical tests, with system checks for MOPS, serial powering, and cooling efficiency. These tests identify issues in integration, module performance, and NTC temperature sensor consistency, ensuring the demonstrator’s readiness for large-scale production.
@@ 3 Readout Chain, DCS, and Data Transmission
The demonstrator system serves as a proof of concept for the ITk design, showcasing and testing the solutions envisioned for the final installation. By minimizing cabling needs through serial powering and using electrical services to route data and command signals to Opto-converters, the demonstrator replicates key aspects of the intended system, providing critical insights into its implementation. Signal aggregation and electrical-to-optical conversion occur in the Opto-box outside the test environment for radiation shielding and component accessibility. Each quad module uplink is supported by 1.28 Gb/s fibers to manage high data volumes, and cable lengths reflect the real setup for bandwidth and transmission reliability testing.
@@
Furthermore, the DCS controls component powers and integrates it into a finite state machine (FSM) interface, simplifying setup operation and safety by the interlock. The FSM allows for quick navigation between component states via a graphical user interface (GUI), enabling efficient operation for the ITk pixel detector’s future deployment.
@@ 4 Local Loaded Supports and $CO_{2}$ Cooling System
Precision in assembly is critical, especially regarding the layer of glue used to attach modules to the stave [10]. Excessive glue can shift module positions and reduce the thermal interface, negatively impacting heat dissipation from modules to the cooling system. Each stave integrates three key elements for optimized thermal management: thin-walled titanium tubes that circulate $CO_{2}$ coolant, base cooling blocks strategically placed to create thermally conductive interfaces, and carbon-fiber composites that form the primary support structure as Figure 3 explicitly shows, ensuring stability while managing temperature fluctuations.
@@
Figure 3: Figure (left) shows $CO_{2}$ flow from inlet to exhaust with the module added to the cell-loading structure on the stave. Figure (right) shows thermal variations in NTC measurements that were observed between single-module cooling jig tests and via the $CO_{2}$ cooling pipe.
@@ 5 RD53A Module Reliability in Full-System Tests
Figure 4: Inclined Half-Ring (IHR) stave installed inside the test box and mounted with 10 cell-loaded RD53A quad modules attached with pig-tail cables via the patch panel 0 (PP0).
A key objective of the OB demonstrator project is to validate the loading concepts crucial to ITk detector construction with the overall sub-systems described in section 3. Using detector prototypes equipped with RD53A electronics [11] as in Figure 4, the electrical scans taken at each stage of the production flow ensure the reliability of the module’s performance. A large set of electrical scans are needed to characterize the functionality of individual module components using standardized quality control (QC) measures across all production stages. To ensure direct inspections of any potential degradation by comparing the output to the previous and following stages. A final round of electrical scans evaluates the fully loaded modules in the ITk OB demonstrator setup, as illustrates 10 modules loaded on the IHR in Figure 5 of the OB demonstrator at SR1 in CERN. Which features the full-system test of DAQ, $CO_{2}$ cooling, SLDO serial powering chains, and collective module testing. The performance across all six testing stages demonstrates consistency, with minimal variation across stages despite environmental differences and fundamentally different setups.
Figure 5: The figure illustrates the multiple testing stages that RD53a modules undergo throughout various production steps. Upon arrival at CERN, an initial test is conducted to evaluate baseline functionality. Then, wire bonding is secured by coating for protection. The modules then proceed to cell loading on the back of the module as shown in Figure 3 and pig-tail assembly before entering the final OB demonstrator system test phase [12]. Notably, noise scan performance remains consistent across all stages, indicating stable behavior for the IHR modules in the OB demonstrator.